OpenStudy (anonymous):

Write an equation in standard form of the line that pages through (7,-3) and has a y-int of 2. Standard form = Ax+By=C

OpenStudy (anonymous):

start with piont slope from and then convert

OpenStudy (anonymous):

I suggest using Slope-Intercept form and then moving over the x term.

OpenStudy (anonymous):

The answer is supposed to be 5x+2y=14 but I have no idea how they got 5x nor 14.

OpenStudy (anonymous):

You have everything needed to solve for the slope using the Slope-Intercept form:\[y=mx+b\]x = 7, y = -3 and b = 2. solve for m. then subtract mx from both sides. -mx + y = 2 if m is a fraction, multiply all terms by the denominator. m = n/d, then -mx + y = b => -nx +dy = db

OpenStudy (anonymous):

do you follow?